I need an equation for the tollowing word problem.
at your lemonade stand you charge .50 for 1/2 cup and .75 for a full cup at the end of the day you see that 12 cups have been used and you have made 8$ how name of each size of drink did you sell?

Answer by mananth(16946)

1/2 cup ---x numbers
full cup---y numbers
x+y =12
0.5x+0.75y=8
multiply by 100
50x+75y=800
these are the two equations solve for x & y
1 x + 1 y = 12 .............1
50 x + 75 y = 800 .............2
Eliminate y
multiply (1)by -75
Multiply (2) by 1
-75 x -75 y = -900
50 x 75 y = 800
Add the two equations
-25 x = -100
/ -25
x = 4 1/2 cup
plug value of x in (1)
1 x + 1 y = 12
4 + 1 y = 12
1 y = 12 -4
1 y = 8
y = 8 full cup